# https://kaist-viclab.github.io/SLiM_site/

### 11 Mar 2026 v1

@@ 0.E.1 Transfer Learning Evaluation
To assess the generalizability of the learned action representations, we conduct a transfer learning evaluation following the established protocol in recent literature. In this scenario, the SLiM encoder $g$ is first pre-trained on a large-scale source dataset using our self-supervised framework. The frozen encoder is then evaluated on a different target dataset using the linear evaluation mechanism where only a single linear classifier is trained on top of the fixed representations without additional fine-tuning. This process verifies whether the model has captured universal motion semantics that are robust to domain shifts and varying skeleton topologies. Table 8 shows the transfer learning performance on PKU-MMD II [18] dataset. As shown, SLiM demonstrates superior transferability across all tested scenarios compared to previous state-of-the-art methods. Specifically, when pre-trained on the NTU-120 [19] X-Sub split, SLiM achieves a peak accuracy of 75.3%, outperforming S-JEPA [1] by 1.1%-point. These results highlight that the representations learned via our Semantic Tube Masking (STM) and Global-Local Contrastive Learning (GLCL) objectives are more context-rich and transferable than those focusing on low-level coordinate reconstruction. By effectively predicting high-dimensional latent features, SLiM establishes a more robust anatomical understanding, ensuring that the captured motion patterns generalize effectively across different datasets and subject distributions.
@@
LongT-GAN [41] 44.8
ISC [29] 51.1 52.3
CMD [23] 56.0 57.0
MacDiff [37] 72.2 73.4
IGM [16] 59.8
SkeletonMAE [38] 58.4 61.0
MAMP [22] 70.6 73.2
S-JEPA [1] 71.4 74.2

### 12 Mar 2026 v2

@@ 0.E.1 Transfer Learning Evaluation
To assess the generalizability of the learned action representations, we conduct a transfer learning evaluation following the established protocol in recent literature. In this scenario, the SLiM encoder $g$ is first pre-trained on a large-scale source dataset using our self-supervised framework. The frozen encoder is then evaluated on a different target dataset using the linear evaluation mechanism where only a single linear classifier is trained on top of the fixed representations without additional fine-tuning. This process verifies whether the model has captured universal motion semantics that are robust to domain shifts and varying skeleton topologies. Table 8 shows the transfer learning performance on PKU-MMD II [PKU] dataset. As shown, SLiM demonstrates superior transferability across all tested scenarios compared to previous state-of-the-art methods. Specifically, when pre-trained on the NTU-120 [NTU120] X-Sub split, SLiM achieves a peak accuracy of 75.3%, outperforming S-JEPA [SJEPA] by 1.1%-point. These results highlight that the representations learned via our Semantic Tube Masking (STM) and Global-Local Contrastive Learning (GLCL) objectives are more context-rich and transferable than those focusing on low-level coordinate reconstruction. By effectively predicting high-dimensional latent features, SLiM establishes a more robust anatomical understanding, ensuring that the captured motion patterns generalize effectively across different datasets and subject distributions.
@@
LongT-GAN [LongTGAN] 44.8
ISC [ISC] 51.1 52.3
CMD [CMD] 56.0 57.0
MacDiff [MacDiff] 72.2 73.4
IGM [IGM] 59.8
SkeletonMAE [SkeletonMAE] 58.4 61.0
MAMP [MAMP] 70.6 73.2
S-JEPA [SJEPA] 71.4 74.2